OpenStudy (anonymous):

What is the value of f(7) for the function f(x) = 4x + 3? (Type only the value of your final answer, no variables.) Numerical Answers Expected!

OpenStudy (ghazi):

put x=7 in the given function f(7)=4*7+3=28+3

OpenStudy (anonymous):

thank you

OpenStudy (ghazi):

:)